Suppose you sell a fixed asset for $96,904 when its book value is $69,541. If your company's marginal tax rate is 25 percent, what will be the after-tax cash flow of this sale? DO NOT USE DOLLAR SIGNS OR COMMAS IN YOUR ANSWER. ROUND ANSWER TO THE NEAREST DOLLAR (example: 18630).
Answer #1

Gain on sale=96,904-69,541=$27363

Hence after-tax cash flow=Sale proceeds-(Gain on sale*Tax rate)

=96,904-(27363*25%)

which is equal to

=90063(Approx).